OpenStudy (anonymous):

Write in set builder notation: A is the set of letters in the word “notations”

OpenStudy (unklerhaukus):

you dont have to repeat terms in a set

OpenStudy (anonymous):

A={xIx=n,o,t,a,i,o,s}

OpenStudy (unklerhaukus):

you made a repeat

OpenStudy (anonymous):

A={xIx=n,o,t,a,i,s}

OpenStudy (anonymous):

Let A=〖10〗^a, B=〖10〗^b , C=〖10〗^c and a+b+c=0 . Find the numerical Value of 〖 A〗^(1/b+1/c) + B^(1/c+1/a)+C^(1/a+1/b) .

OpenStudy (anonymous):

Let a, b, c be positive integers. x, y, z be real numbers and a<b<c. If a^x= b^y=c^z=〖30〗^w and 1/x+ 1/y+ 1/z= 1/w . Find the numerical values of a, b, and c.

OpenStudy (raden):

new question u better post to (ask a question place) so that every one can see it